# Perturbation-Restrained Sequential Model Editing

### v1

@@ 4 PRUNE: Perturbation Restraint on Upper bouNd for Editing
According to the analysis in Section 3, the bottleneck of the general abilities during sequential editing lies in the escalating value of the condition number. In this section, a framework termed Perturbation Restraint on Upper bouNd for Editing (PRUNE) is proposed, which applies the condition number restraints to preserve general abilities and maintain new editing knowledge simultaneously.
@@
Given an edited matrix with $N$ edits, $W_{N}=W+\sum_{j=1}^{N}\Delta W_{j}$, as shown in Figure 2, its maximum singular value is constantly increasing, while the minimum singular value is basically unchanged as the number of edits $N$ increases. This directly leads to the increasing condition number of the edited matrix. Therefore, our motivation is to restrain the maximum singular value of the edited matrix to lower the upper bound on the perturbation. If we directly perform SVD operation on $W_{N}$ and reduce its singular values, the original $W$ will be inevitably destroyed. Consequently, an analysis of the singular values of $\sum_{j=1}^{N}\Delta W_{j}$ is conducted, and the results in Table 1 present that its maximum singular value becomes very large when $N$ is large. We assume that this is the main reason why the maximum singular value of $W_{N}$ is large, our method therefore aims to restrain the large singular values of $\sum_{j=1}^{N}\Delta W_{j}$.
@@
In this way, the condition number of the edited matrix is reduced (see Appendix C.4) and the upper bound on perturbation is significantly restrained.
@@ 5.1 Base LLMs and Editing Methods
Experiments were conducted on three LLMs including GPT-2 XL (1.5B) [48], LLaMA-2 (7B) [53] and LLaMA-3 (8B) [^4] https://llama.meta.com/llama3/. Three popular editing methods were selected as the baselines including MEND [43], ROME [41], and MEMIT [42]. Readers can refer to Appendix B.1 for the details of these editing methods.
@@ 5.2 Editing Datasets and Evaluation Metrics
Two popular model editing datasets Zero-Shot Relation Extraction (ZsRE) [33] and CounterFact [41] were adopted in our experiments. ZsRE is a QA dataset using question rephrasings generated by back-translation as the equivalence neighborhood. A key distinction between CounterFact and ZsRE datasets is that ZsRE contains true facts, while CounterFact contains counterfactual examples where the new target has a lower probability when compared to the original answer [22]. Readers can refer to Appendix B.2 for examples of each dataset.
To assess the editing performance of editing methods, following previous works, three fundamental metrics were employed: efficacy, generalization and locality [3, 43, 41, 42]. Given an original model $f_{\theta_{0}}$, an edited model $f_{\theta_{n}}$ with $n$ times sequential editing. Each edit $e_{i}=(s_{i},r_{i},o_{i},o_{i}^{*})$ has an editing prompt $p_{i}$, paraphrase prompts $\mathcal{P}^{G}_{i}$, and locality prompts $\mathcal{P}^{L}_{i}$.
Different from previous studies that assess the edited models after each individual edit [22, 62], this paper evaluated whether the final edited models after completing all edits can still recall all preceding edits, which is more challenging and common in real-world.
@@ 5.3 Downstream Tasks, Datasets and Metrics
To explore the side effects of sequential model editing on the general abilities of LLMs, four representative tasks with corresponding datasets were adopted for assessment, including:
Reasoning on the GSM8K [8], and the results were measured by solve rate.
Summarization on the SAMSum [19], and the results were measured by the average of ROUGE-1, ROUGE-2 and ROUGE-L following Lin [36].
Open-domain QA on the Natural Question [31], and the results were measured by exact match (EM) with the reference answer after minor normalization as in Chen et al. [4] and Lee et al. [32].
Natural language inference (NLI) on the RTE [11], and the results were measured by accuracy of two-way classification.
For each dataset, some examples were randomly sampled for evaluation. Details of prompts for each tasks were shown in Appendix B.4.
Figure 3: The downstream task performance (%) of models edited by three editing methods with LLaMA-2 (7B) on the CounterFact dataset. The dashed lines refer to the results of the unrestrained editing methods. The solid lines refer to the results of the editing methods coupled with the proposed PRUNE framework. Statistical significance tests were performed to demonstrate that the improvement in PRUNE compared to baseline was statistically significant (t-test with $p$-value <$0.05$).
5.4 Results of General Abilities
Figure 3 illustrates the downstream task performance of edited models with LLaMA-2 (7B) on the CounterFact dataset. Due to page limitation, results of other LLMs and datasets were put in Appendix C.1. These results were analyzed from the following perspectives.
Current editing methods significantly compromised general abilities. As depicted by the dashed lines of Figure 3, both the ROME and MEMIT methods initially maintained relatively stable performance in downstream tasks when the number of edits was small ($\leq$ 100). However, as the number of edits surpassed 200, a noticeable decline in performance was observed across all tasks for both methods. Additionally, the MEND method exhibited significant performance degradation after just 20 sequential edits, indicating its inadequacy as a sequential model editing method. Furthermore, when comparing LLMs of different sizes, a general trend emerged: larger models suffered more pronounced compromises in their general abilities when subjected to the same number of edits. For instance, with 300 edits, MEMIT’s performance on GPT2-XL remained largely unchanged, whereas it dwindled to nearly 0 on LLaMA-2 and LLaMA-3.
@@ 5.4 Results of General Abilities
The proposed PRUNE can preserve considerable general abilities. As shown by the solid lines of Figure 3, when MEMIT was coupled with PRUNE and subjected to 200 edits, its downstream tasks performance remained close to that of the unedited model. However, for the unrestrained MEMIT, downstream task performance had plummeted to nearly 0 by this point. This consistent trend was also observed with ROME and MEND. Nevertheless, for models edited using the unrestrained MEND method, performance degradation was stark after just 20 edits. Even with the addition of PRUNE, preservation could only be extended up to 40 edits. This suggests that while PRUNE effectively preserves general abilities, it does have an upper limit determined by the unrestrained editing method.
Figure 4: The editing performance (%) of three editing methods with LLaMA-2 (7B) on the CounterFact dataset. The dashed lines refer to the results of the unrestrained editing methods. The solid lines refer to the results of the editing methods coupled with the proposed PRUNE. Statistical significance tests were performed to demonstrate that the improvement in PRUNE compared to baseline was statistically significant (t-test with $p$-value <$0.05$).
5.5 Results of Editing Performance
Figure 4 shows different metrics used for measuring the editing performance of edited models with LLaMA-2 (7B) on the CounterFact dataset. Other results across models and datasets were put in Appendix C.2. Three conclusions can be drawn here.
Previous editing facts were forgotten as the number of edits increased. As shown by the dashed lines of Figure 4, the decline in efficacy and generalization suggests that in sequential editing scenarios, post-edited models gradually forget knowledge acquired from previous edits after a few iterations. Comparing these editing methods, we also observed a notable drop in efficacy and generalization after hundreds of edits with ROME and MEMIT, whereas these values decreased significantly after only 30 edits with MEND. This indicates that in sequential editing scenarios, the MEND method struggled to successfully integrate new knowledge into LLMs after several edits.
@@ 5.5 Results of Editing Performance
PRUNE can effectively maintain the editing performance. This is shown by the solid lines of Figure 4 and could be analyzed from two aspects. On the one hand, when the number of edits was small, the editing performance of each editing method coupled with PRUNE was about the same as the unrestrained method. On the other hand, it significantly mitigated the forgetting of editing facts and the perturbation of irrelevant facts when the number of edits was large during the sequential editing. Specifically, when the number of edits reached 150, the editing performance of MEMIT was very low. But when coupled with PRUNE, its performance remained relatively stable.
5.6 Editing Facts Forgetting Analysis
@@
Figure 5: 2-dimensional PCA visualization of first 100 values. The model was edited by ROME with LLaMA-3 (8B) on the CounterFact dataset.
@@
In section 3, the analysis was conducted to elucidate the reasons behind the degradation in general abilities with an increasing number of edits. Subsequent experiments quantitatively demonstrated the effectiveness of the proposed PRUNE. Here, we delve into qualitative analysis to explain why editing facts are forgotten and how PRUNE can mitigate this forgetting.
@@
Initially, a set of editing facts =$\mathcal{E}=\left\{e_{1},e_{2},\ldots\right\}$ was collected, where $\mid\mathcal{E}\mid=200$. ROME was employed for analysis, and the original matrix was defined as $W$. During sequential editing, ROME computed key-value pairs $(k^{e}_{j},v^{e}_{j})$ of the last subject token to generate $\Delta W_{j}$ for each edit $e_{j}$ to incorporate new facts, satisfying the equation: $W_{j}\cdot k^{e}_{j}=v^{e}_{j}$. However, when evaluating editing performance, the edited model obtained from the last edit was utilized, thus computing values [^5] Since ROME only modifies one matrix, the $k^{e}_{j}$ remains the same across these edited models.: $W_{200}\cdot k^{e}_{j}=\hat{v}^{e}_{j}$. After adopting PRUNE to ROME, this equation became $\overline{W}_{200}\cdot k^{e}_{j}=\overline{v}^{e}_{j}$. We hypothesized that if $\hat{v}^{e}_{j}$ was similar to $v^{e}_{j}$, the editing fact $e_{j}$ could be maintained.
@@
Denote $V_{Current}=\{v^{e}_{j}\}$, $V_{Editing}=\{\hat{v}^{e}_{j}\}$, and $V_{Prune}=\{\overline{v}^{e}_{j}\}$. Specifically, these corresponding values of the first 100 edits were used, as they are more prone to be forgotten than the last 100. Principal Component Analysis (PCA) [18] was employed to visualize these values. The first two principal components of each value were calculated and illustrated, as they can represent most of its features [66]. As shown in Figure 5, on the one hand, the discrepancy between the principal components of $V_{Current}$ and $V_{Editing}$ was markedly large. This indicates that after 200 edits to the model, the values corresponding to the first 100 facts stored in the edited matrix are severely corrupted, leading to significant forgetfulness. On the other hand, after adopting PRUNE, the discrepancy between the principal components of $V_{Current}$ and $V_{Prune}$ was small. This demonstrates that PRUNE effectively maintains the values and mitigates the forgetting of editing facts.
6 Conclusion
@@
Limitations & Future Work
The limitations of our work are discussed as follows. Firstly, while this paper focuses on editing a single fact at a time in sequential model editing, some studies have explored updating hundreds or thousands of facts simultaneously in batch editing. Therefore, investigating batch-sequential editing could enhance the scalability of model editing and remains further research. Secondly, for the experimental settings, it is necessary to explore the performance of larger-size models and more editing methods on more downstream tasks. Thirdly, the current focus of editing knowledge primarily revolves around factual knowledge. However, it is also important to investigate whether editing other types of knowledge will affect general abilities and whether PRUNE is effective in this situation. Additionally, the proposed PRUNE is only applied once after the completion of the last edit. But it could also be utilized multiple times during the sequential editing process, and we intuitively believe that this approach would be more conducive to preserving the general abilities of the model. These aspects are yet to be fully understood and warrant a more comprehensive study.

### v3

@@ 4 PRUNE: Perturbation Restraint on Upper bouNd for Editing
Motivation
According to the analysis in Section 3, the bottleneck of the general abilities during sequential editing lies in the escalating value of the condition number. Assuming a set of edits $\{e_{i}\}$ and their corresponding edit update matrices $\{\Delta W_{i}\}$, the information contained in these edit update matrices coordinates with each other to a certain extent since the parametric knowledge of LLMs is distributional rather than independent. This editing overfitting is reflected in SVD, where the largest singular value of the edited matrix $W_{N}$ becomes significantly large after the addition of these edit update matrices. To illustrate this, consider an extreme example: suppose we make $N$ edits, where each edit changes the answer to the question “Who is the president of the United States?” to “Biden”. Each edit update matrix is denoted as $\Delta W_{1}$, and its maximum singular value is $\delta_{max}$. Then the sum of the $N$ edit update matrices is $N\Delta W_{1}$, and its maximum singular value is $N\delta_{max}$, which is amplified by $N$ times. Therefore, our goal is to reduce the editing overfitting in edited matrix $W_{N}$ as much as possible while also retaining valuable editing information. In this section, a framework termed Perturbation Restraint on Upper bouNd for Editing (PRUNE) is proposed, which applies the condition number restraints to preserve general abilities and maintain new editing knowledge.
@@
Given an edited matrix with $N$ edits, $W_{N}=W+\sum_{j=1}^{N}\Delta W_{j}$, as shown in Figure 2, its maximum singular value is constantly increasing, while the minimum singular value is basically unchanged as the number of edits $N$ increases. This directly leads to the increasing condition number of the edited matrix. Therefore, our motivation is to restrain the large singular value of the edited matrix to lower the upper bound on the perturbation. If we directly perform SVD operation on $W_{N}$ and reduce its singular values, the original $W$ will be inevitably destroyed. Consequently, an analysis of the singular values of $\sum_{j=1}^{N}\Delta W_{j}$ is conducted, and the results in Table 1 present that its maximum singular value becomes very large when $N$ is large. Since the singular values of $W$ are relatively small, we can assume that the large maximum singular value of $\sum_{j=1}^{N}\Delta W_{j}$ is the main reason why the maximum singular value of $W_{N}$ is large, our method therefore aims to restrain the large singular values of $\sum_{j=1}^{N}\Delta W_{j}$.
@@
In this way, the condition number of the edited matrix is reduced (see Appendix C.4) and the upper bound on perturbation is significantly restrained. It is worth noting that the PRUNE proposed here is only used once in Section 5, but can actually be used multiple times in the editing process to better maintain the general ability. We provide a comparison of the two strategies in Appendix C.6.
@@ 5.1 Base LLMs and Editing Methods
Experiments were conducted on three LLMs including GPT-2 XL (1.5B) (Radford et al., 2019), LLaMA-2 (7B) (Touvron et al., 2023) and LLaMA-3 (8B) [^4] https://llama.meta.com/llama3/. Three popular editing methods were selected as the baselines including MEND (Mitchell et al., 2022a), ROME (Meng et al., 2022), and MEMIT (Meng et al., 2023). Appendix B.1 shows the details of these editing methods.
@@ 5.2 Editing Datasets and Evaluation Metrics
To make a more comprehensive evaluation, we used two types of knowledge for editing: factual knowledge and conceptual knowledge. (1) For factual knowledge, two popular model editing datasets Zero-Shot Relation Extraction (ZsRE) (Levy et al., 2017) and CounterFact (Meng et al., 2022) were adopted in our experiments. These two datasets are QA datasets. A key distinction between CounterFact and ZsRE datasets is that ZsRE contains true facts, while CounterFact contains counterfactual examples where the new target has a lower probability when compared to the original answer (Gupta et al., 2024b). (2) For conceptual knowledge, the ConceptEdit dataset (Wang et al., 2024) was adopted. Due to the limitations of computing resources and pages, most of the experiments in this paper were conducted on factual datasets, with the results presented in Sections 5.4 and 5.5. Meanwhile, Section 5.6 provided some results on conceptual datasets. Readers can refer to Appendix B.2 for examples of each dataset.
To assess the editing performance of editing methods, following previous works (Cao et al., 2021; Mitchell et al., 2022a; Meng et al., 2022; 2023; Ma et al., ), three fundamental metrics were employed: efficacy, generalization and locality. Given an original model $f_{\theta_{0}}$, an edited model $f_{\theta_{n}}$ with $n$ times sequential editing. Define $\mathbbm{1}$ as the indicator function. Each edit $e_{i}=(s_{i},r_{i},o_{i},o_{i}^{*})$ has an editing prompt $p_{i}$, paraphrase prompts $\mathcal{P}^{G}_{i}$, and locality prompts $\mathcal{P}^{L}_{i}$.
Different from previous studies that assess the edited models after each individual edit (Gupta et al., 2024b; Yao et al., 2023), this paper evaluated whether the final edited models after completing all edits can still recall all preceding edits, which is more challenging and common in real-world.
@@ 5.3 Downstream Tasks, Datasets and Metrics
To explore the side effects of sequential model editing on the general abilities of LLMs, four representative tasks with corresponding datasets were adopted for assessment following previous work (Gu et al., 2024; Gupta et al., 2024b; Lin et al., 2024; Zhang et al., 2024), including:
Reasoning on the GSM8K (Cobbe et al., 2021), and the results were measured by solve rate.
Summarization on the SAMSum (Gliwa et al., 2019), and the results were measured by the average of ROUGE-1, ROUGE-2 and ROUGE-L following Lin (2004).
Open-domain QA on the Natural Question (Kwiatkowski et al., 2019), and the results were measured by exact match (EM) with the reference answer after minor normalization as in Chen et al. (2017) and Lee et al. (2019).
Natural language inference (NLI) on the RTE (Dagan et al., 2005), and the results were measured by accuracy of two-way classification.
For each dataset, some examples were randomly sampled for evaluation. Details of prompts for each task were shown in Appendix B.4.
Figure 3: The downstream task performance (%) of models edited by three editing methods with LLaMA-2 (7B) on the ZsRE dataset. The dashed lines refer to the results of the unrestrained editing methods. The solid lines refer to the results of the editing methods coupled with the proposed PRUNE framework. Statistical significance tests were performed to demonstrate that the improvement in PRUNE compared to baseline was statistically significant (t-test with $p$-value <$0.05$).
5.4 General Abilities Results on Factual Knowledge
Figure 3 illustrates the downstream task performance of editing methods with LLaMA-2 (7B) on the ZsRE dataset. Due to page limitation, results of other LLMs and factual datasets were put in Appendix C.1. These results were analyzed from the following perspectives.
Current editing methods significantly compromised general abilities. As depicted by the dashed lines of Figure 3, both the ROME and MEMIT methods initially maintained relatively stable performance in downstream tasks when the number of edits was small ($\leq$ 50). However, as the number of edits surpassed 100, a noticeable decline in performance was observed across all tasks for both methods. Additionally, the MEND method exhibited significant performance degradation after just 20 sequential edits, indicating its inadequacy as a sequential model editing method. Furthermore, when comparing LLMs of different sizes, a general trend emerged: larger models suffered more pronounced compromises in their general abilities when subjected to the same number of edits. For instance, with 300 edits, MEMIT’s performance on GPT2-XL remained largely unchanged, whereas it dwindled to nearly 0 on LLaMA-2 and LLaMA-3.
@@ 5.4 General Abilities Results on Factual Knowledge
The proposed PRUNE can preserve considerable general abilities. As shown by the solid lines of Figure 3, when MEMIT was coupled with PRUNE and subjected to 100 edits, its downstream tasks performance remained close to that of the unedited model. However, for the unrestrained MEMIT, downstream task performance had plummeted to nearly 0 by this point. This consistent trend was also observed with ROME and MEND. Nevertheless, for models edited using the unrestrained MEND method, performance degradation was stark after just 10 edits. Even with the addition of PRUNE, preservation could only be extended up to 20 edits. This suggests that while PRUNE effectively preserves general abilities, it does have an upper limit determined by the unrestrained editing method.
Figure 4: The editing performance (%) of editing methods with LLaMA-2 (7B) on the ZsRE dataset. The dashed lines refer to the results of the unrestrained editing methods. The solid lines refer to the results of the editing methods coupled with the proposed PRUNE. Statistical significance tests were performed to demonstrate that the improvement in PRUNE compared to baseline was statistically significant (t-test with $p$-value <$0.05$).
5.5 Editing Performance Results on Factual Knowledge
Figure 4 shows three metrics used for measuring the editing performance with LLaMA-2 (7B) on the ZsRE dataset. Other results were put in Appendix C.2. Three conclusions can be drawn.
Previous editing facts were forgotten as the number of edits increased. As shown by the dashed lines of Figure 4, the decline in efficacy and generalization suggests that in sequential editing scenarios, post-edited models gradually forget knowledge acquired from previous edits after a few iterations. Comparing these editing methods, we also observed a notable drop in efficacy and generalization after hundreds of edits with ROME and MEMIT, whereas these values decreased significantly after only 15 edits with MEND. This indicates that in sequential editing scenarios, the MEND method struggled to successfully integrate new knowledge into LLMs after several edits.
@@ 5.5 Editing Performance Results on Factual Knowledge
PRUNE can effectively maintain the editing performance. This is shown by the solid lines of Figure 4 and could be analyzed from two aspects. On the one hand, when the number of edits was small, the editing performance of each editing method coupled with PRUNE was about the same as the unrestrained method. On the other hand, it significantly mitigated the forgetting of editing facts and the perturbation of irrelevant facts when the number of edits was large during the sequential editing. Specifically, when the number of edits reached 100, the editing performance of MEMIT was very low. But when coupled with PRUNE, its performance remained relatively stable. These observations further validate our motivation in Section 4, demonstrating that the information in the edit update matrices is coordinated, and that performing too many edits can easily result in overfitting. Therefore, applying a certain degree of restraint to edit perturbations can help preserve the model’s general abilities while maintaining the editing knowledge.
5.6 Editing with Conceptual Knowledge
Section 5.4 and 5.5 analyzed the results on factual knowledge. This section conducted some experiments with ROME on conceptual knowledge using the ConceptEdit dataset (Wang et al., 2024) to make a more comprehensive evaluation. For editing performance, in addition to the three basic metrics, this dataset also designed a new metric “Instance Change” to measure whether the instances under the concept changed accordingly when the definition of the concept was changed.
As shown in Table 6, the performance trends of editing and downstream tasks were similar to those observed with the factual datasets. But there are several key differences: (1) When the number of edits was the same, the editing performance of conceptual knowledge was lower than that of factual knowledge. (2) Both editing performance and general abilities deteriorated more quickly than factual knowledge. For example, even if the number of edits was 100, the editing performance and downstream task performance of ROME were very low, while it was still relatively high when editing factual knowledge. (3) The low “Instance Change” indicated that when the definition of a concept was altered, the instances contained in the original concept were still recognized by the model as belonging to that concept. This shows that this editing method primarily modifies the definition without successfully altering the relationship between concepts and instances, which is not reasonable. These findings indicate that conceptual knowledge is more abstract and more difficult to edit than factual knowledge, highlighting the need to explore editing methods for different types of knowledge.
Table 2: Evaluation results (%) of LLaMA-2 (7B) edited by ROME on the ConceptEdit dataset.
Mode General Abilities Editing Performance
Method Edits Reasoning Summa Open-QA NLI Efficacy General Locality Instance
ROME 20 75.13 11 6.50 24.7 49.15 52.58 35.68 25
50 20.67 4.90 1.50 0.7 55.42 49.45 19.94 12
100 12.29 4.7 0.77 0 28.25 30.18 5.68 10
200 0 4.62 0 0 10.14 8.65 5.31 -8.99
ROME+PRUNE 20 89.38 14.34 23.37 63.54 75.66 58.35 71.7 25
50 85.15 14.06 25.29 50.52 56.51 45.55 73.16 8
100 90.78 13.75 21.46 53.17 46.22 42.26 64.06 20
200 72.9 10.55 22.22 46.15 35.82 34.95 46.65 32
5.7 Analysis on the Forgetting of Editing Facts
@@
Figure 5: 2-dimensional PCA visualization of first 100 values. The model was edited by ROME with LLaMA-2.
@@
Section 3 conducted analysis to elucidate the reasons behind the degradation in general abilities with an increasing number of edits. Subsequent experiments quantitatively demonstrated the effectiveness of PRUNE. Here, we delve into qualitative analysis to explain why editing facts are forgotten and how PRUNE can mitigate this forgetting.
@@
Initially, given a set of editing facts =$\mathcal{E}=\left\{e_{1},e_{2},\ldots\right\}$, where $\mid\mathcal{E}\mid=200$. ROME was employed for analysis, and the original matrix was defined as $W$. During sequential editing, ROME computed key-value pairs $(k^{e}_{j},v^{e}_{j})$ of the last subject token to generate $\Delta W_{j}$ for each edit $e_{j}$ to incorporate new facts, satisfying the equation: $W_{j}\cdot k^{e}_{j}=v^{e}_{j}$. However, when evaluating editing performance, the edited model obtained from the last edit was utilized, thus computing values [^5] Since ROME only modifies one matrix, the $k^{e}_{j}$ remains the same across these edited models.: $W_{200}\cdot k^{e}_{j}=\hat{v}^{e}_{j}$. After adopting PRUNE to ROME, this equation became $\overline{W}_{200}\cdot k^{e}_{j}=\overline{v}^{e}_{j}$. We hypothesized that if $\hat{v}^{e}_{j}$ was similar to $v^{e}_{j}$, the editing fact $e_{j}$ could be maintained.
@@
Denote $V_{Current}=\{v^{e}_{j}\}$, $V_{Editing}=\{\hat{v}^{e}_{j}\}$, and $V_{Prune}=\{\overline{v}^{e}_{j}\}$. Specifically, these corresponding values of the first 100 edits were used, as they are more prone to be forgotten than the last 100. Principal Component Analysis (PCA) (Gewers et al., 2022) was employed to visualize these values. The first two principal components of each value were calculated and illustrated, as they can represent most of its features (Zheng et al., ). As shown in Figure 5, on the one hand, the discrepancy between the principal components of $V_{Current}$ and $V_{Editing}$ was markedly large. This indicates that after 200 edits to the model, the values corresponding to the first 100 facts stored in the edited matrix are severely corrupted, leading to significant forgetfulness. On the other hand, after adopting PRUNE, the discrepancy between the principal components of $V_{Current}$ and $V_{Prune}$ was small. This demonstrates that PRUNE effectively maintains the values and mitigates the forgetting of editing facts.
6 Conclusion and Limitation
@@
Limitation Firstly, this paper focuses on editing a single fact at a time in sequential model editing, but some works study updating hundreds of facts simultaneously in batch editing. Therefore, investigating batch-sequential editing could enhance the scalability of model editing. Secondly, it is necessary to explore the performance of larger-size models and more editing methods on more downstream tasks.
Acknowledgments
We would like to express gratitude to the anonymous reviewers for kind comments. This work is funded by the National Science and Technology Major Project (No. 2023ZD0121103).
Appendix